# The Fourier spectral approach to the spatial discretization of quasilinear hyperbolic systems

### v1

@@ 4.2 Numerical experiments in dimension one
where $\hat{\bm{\eta}}=(\hat{\bm{\eta}}_{-M+1},\ldots,\hat{\bm{\eta}}_{M}),\hat{\bm{u}}=(\hat{\bm{u}}_{-M+1},\ldots,\hat{\bm{u}}_{M})$ are the coefficients of the Fast Fourier transform of $\bm{\eta},\bm{u}$, and $\hat{k}=k/\pi,k=-M+1,\ldots,M$ are the discrete Fourier modes. Abusing notation, we will incorrectly refer to $\hat{\bm{\eta}}_{k},\hat{\bm{u}}_{k}$ as Fourier coefficients (they are related to the coefficients $c_{j}$ of infinite Fourier series $f(x)=\sum_{j\in\mathbb{Z}}c_{j}\exp({\rm i}jx)$ through $\hat{\bm{f}}_{k}=\sum_{j\in\mathbb{Z}}c_{k+2jM}$). For functions $f\in H^{s}(2\pi\mathbb{T})$, the error due to this aliasing effect is of order $M^{-s}$.
Spatial differentiation is now obtained by multiplying the Fourier coefficients with ${\rm i}k$. Nonlinear operations are computed pointwise on collocation points $x_{n}$, via inverse Fast Fourier transform. This procedure leads, in general, to aliasing errors. For polynomial nonlinearities (such as for the Saint-Venant system), one can use so-called dealiasing techniques to remove these errors. For quadratic nonlinearities, one may for example use Orszag’s $3/2$-rule [20], which consists of adding a sufficient number of Fourier modes with coefficients set to zero. For more information on spectral methods and dealiasing techniques, we refer to [5] and [25].
@@
From now on, we denote the number of collocation points by $2M$, and let $N=\lfloor 2M/3\rfloor$, that is, the greatest integer smaller than $32M/$ 3. Our numerical scheme maintains the highest $1/3$ of the Fourier modes to zero at each time-step, see the discussion above. Abusing notation, we will refer to the fully-discretized numerical solution as $\bm{U}_{N}=(\eta_{N},u_{N})$, since only $2N$ Fourier modes are nonzero. This convention means that $N$ plays the same role in this section as in the previous, analytical sections. We will compute the numerical solution with $152M=2^{j},j=6,\ldots,$ 15 collocation points and use time step $dt=1e-5$. The time step is an order of magnitude smaller than needed to avoid stability issues, and small enough to ensure the error due to the spatial discretization dominates. We will use the solution computed with $2M=2^{15}$ and sharp low-pass filter as a reference solution $\bm{U}_{\rm ref}=(\eta_{\rm ref},u_{\rm ref})$, and compute the relative error of the numerical solutions $\bm{U}_{N}=(\eta_{N},u_{N})$ by comparing with the reference solution:
@@
$\eta_{0}(x)=\frac{1}{2}\exp(-\left|x\right|^{\alpha})\exp(-4x^{2}),\qquad u_{0}(x)=0.$ (4.9)
Notice $\bm{U}^{0}=(\eta_{0},u_{0})$ satisfies both $01+\eta_{0}>$ 0 and the stricter condition $01+\eta_{0}-(u_{0})^{2}>$ 0. The initial surface is a heap of water situated at the origin. Both $\eta_{0},u_{0}$ decay to machine precision at $-\pi,\pi$ and can therefore be seen as periodic. Moreover, $\bm{U}^{0}\in H^{\alpha+1/2}(2\pi\mathbb{T})^{2}$. We let $\alpha=1.5$ and simulate the time-evolution up to a final time $T=0.5$ with either sharp or smooth lowpass filters applied to the nonlinear terms. A plot of the initial data as well as the decay of its Fourier coefficients (through $E^{s}(\bm{U}^{0})=\frac{\left|\bm{U}_{\rm ref}-{\sf P}_{M}\bm{U}_{\rm ref}\right|_{H^{s}}}{\left|\bm{U}_{\rm ref}\right|_{H^{s}}}$) is shown in Figure 1.
@@
Figure 2(a) show log-log plots of the error $E_{s}$ at time $T=0.5$ for the numerical solution for $152M=2^{j},j=6,\ldots,$ 15 computed using sharp and smooth low-pass filters. Figure 2(a) shows the relative error measured in the $L^{2}$-norm, $E_{0}(\bm{U}_{N})$ and in the $H^{1}$-norm, $E_{1}(\bm{U}_{N})$.
@@
$EOC_{s}=\frac{\log(E_{s}(\bm{U}_{N})/E_{s}(\bm{U}_{N+1}))}{\log(2)},$
is given in Figure 2(b). With the initial data in $H^{2}(2\pi\mathbb{T})^{d}$, we expect from the analysis that the $L^{2}$-error should decay as ${\mathcal{O}}(N^{-2})$ and the $H^{1}$-error should decay as ${\mathcal{O}}(N^{-1})$ when using both the sharp and smooth low-pass filter. This aligns with our numerical results. The absolute error is slightly larger when using the smooth low-pass filter, which is to be expected since applying ${\sf S}_{N}$ removes more information in each time step than does ${\sf P}_{N}$. Taking as initial data (4.9) with other values of $\alpha$ (we have tested $\alpha=1,2.5,3$) also yields the expected results.
Let us now consider initial data $(\eta_{0},u_{0})\in{\mathcal{U}}\setminus{\mathcal{U}}_{\mathscr{H}}$, that is satisfying $01+\eta_{0}>$ 0, but not the stricter condition $01+\eta_{0}-(u_{0})^{2}>$ 0. In particular, we will consider the initial data from the example at the end of Section 4.1, with $q=0,p=1,s=2$:
$\bm{U}_{N}^{0}=(\eta_{0},u_{0,N}),\quad\eta_{0}^{N}(x)=-\frac{1}{2}\cos(x),\quad u_{0,N}(x)=\sin(x)+\frac{\sin(Nx)}{N^{2}}$ (4.10)
Notice $f(x)=1-\frac{1}{2}\cos(x)-\sin^{2}(x)<0$ for $x\in(-\frac{\pi}{2},-\frac{\pi}{3})\cup(\frac{\pi}{3},\frac{\pi}{2})$. On the other hand, $01+\eta_{0}(x)>$ 0 for all $x\in 2\pi\mathbb{T}$. We therefore expect the numerical scheme to converge when using the smooth low-pass filter, but that there may be instabilities when discretizing with the sharp low-pass filter. Convergence plots for the numerical solutions in both cases are shown in Figure 3(b). In line with the analysis, we have convergence when using the smooth low-pass filter. However, this is also true for the sharp low-pass filter. Despite many attempts, we have not been able to observe numerical instabilities for initial data we have tested satisfying $01+\eta_{0}>$ 0 but violating $01+\eta_{0}-(u_{0})^{2}>$ 0.
@@
Interestingly, we do observe a difference between sharp and smooth low-pass filters when (barely) violating the non-cavitation assumption $01+\eta_{0}>$ 0. Let
@@
Figure 4 shows the second derivative of the velocity at time $T=0.1$ for smooth and sharp low-pass filters for $2M=2^{10}$ and $2M=2^{12}$. The second derivative blows up at the point $x=0$ where the non-cavitation assumption is violated when using the sharp low-pass filter, but not when using the smooth low-pass filter.
@@ 4.3 Numerical experiments in dimension two
The numerical simulations of the Saint-Venant system with $d=2$ are analogous to the case when $d=1$ and are made in the same Julia package WaterWaves1D [9]. They are also reproducible using the scripts available at WaterWaves1D.jl/examples/StudySaintVenant.jl. We let $2M$ denote the number of collocation points in each of the two spatial dimensions, $x$ and $y$, which form a grid with $4M^{2}$ collocation points. As in the previous section, we set $N=\lfloor 2M/3\rfloor$ and let ${\sf P}_{N},{\sf S}_{N}$ be as described in the introduction. In particular, notice ${\sf S}_{N}$ is now a composition of one dimensional low-pass filters: ${\sf S}_{N}\coloneq\operatorname{Diag}(S_{N}(D))$ with $S_{N}(k_{1},k_{2})\coloneq S_{1}(k_{1}/N)S_{1}(k_{2}/N)$, where we set $S_{1}(\cdot)\coloneq\max\big{(}0,\min\big{(}1,2-2|\cdot|\big{)}\big{)}^{2}$. We denote numerical approximations by $\eta_{N},u_{N},v_{N}$. The values at collocation points $\bm{\eta},\bm{u},\bm{v}$ and associated discrete Fourier modes ^$\bm{\hat{}}{\bm{\eta}},\hat{\bm{u}},\hat{\bm{v}}$ are $2M$-by-$2M$ matrices.
Our main interest in studying numerically the systems in two dimensions is to examine whether we observe any difference between the Hamiltonian and non-Hamiltonian version of the Saint-Venant system, respectively (4.2) and (4.1). While there is indeed a difference with respect to stability of the numerical schemes —we observe instabilities with the sharp low-pass filter when violating $01+\eta_{0}-(u_{0}^{2}+v_{0}^{2})>$ 0 for (4.2), but not for (4.1)— this can be explained by the difference between the hyperbolicity domains ${\mathcal{U}}_{\mathscr{H}}$ and ${\mathcal{U}}$ rather than on the presence or absence of a Hamiltonian structure. Because the setting of dimension $d=2$ is computationally costlier, we use timestep $dt=5e-4$ in our numerical experiments and, when calculating relative errors, we take as a reference solution the numerical solution computed with $2M=2^{10}$.
@@
$\begin{split}\eta_{0}(x,y)&=(h_{0}-1)\cos(x)\cos(y),\\ u_{0}(x,y)&=u_{{\rm l}}\sin(x)\cos(y)+u_{{\rm h}}\frac{\sin(Nx)\cos(Ny)}{N^{s}},\\ v_{0}(x,y)&=v_{{\rm l}}\cos(x)\sin(y)+v_{{\rm h}}\frac{\cos(Nx)\sin(Ny)}{N^{s}},\end{split}$ (4.12)
@@
For the standard, non-Hamiltonian Saint-Venant system (4.1) our numerical results when $d=2$ align with the numerical results when $d=1$. That is, the numerical approximation converges with order $s$ for tested initial data in $H^{s}(2\pi\mathbb{T})$ as long as $01+\eta_{0}>$ 0 when using both the smooth and sharp low-pass filters. Figure 5(a) shows a log-log plot of the relative error measured in the $L^{2}$-norm, $E_{0}(\bm{U}_{N})$ and in the $H^{1}$-norm, $E_{1}(\bm{U}_{N})$, at time $T=0.1$ for $92M=2^{j},j=5,\ldots,$ 9 and $N=\lfloor 2M/3\rfloor$ with initial data (4.12) with $h_{0}=0.5,u_{{\rm l}}=-v_{{\rm l}}=0.5,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$. Analogous results also hold for other values of $s$ (we have tested $s=2.5,3$).
@@
Figure 5: Plot illustrating the convergence of the numerical schemes (4.8) and (4.7) for the systems (4.1) (in the left) and (4.2) (in the right) in two spatial dimensions as the number of collocation points $2M$ increases. The plot shows the relative error of the numerical solution for initial data (4.12) with $h_{0}=0.5,u_{{\rm l}}=-v_{{\rm l}}=0.5,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$ at time $T=0.1$. The initial data is in in $H^{2}((2\pi\mathbb{T})^{2})^{3}$ and the relative error is measured in the $L^{2}$-norm, $E_{0}$ and in the $H^{1}$-norm, $E_{1}$ for $92M=2^{j},j=6,\ldots,$ 9 when using either sharp or smooth low-pass filters. To illustrate, the blue and orange lines have slopes $-2$ and $-1$ respectively. The numerical scheme exhibits spectral convergence with both sharp and smooth low-pass filters, for both systems.
@@
We have not observed any instabilities in the numerical approximation of system (4.1) due to the use of sharp low-pass filter for any of the initial data we tested satisfying $01+\eta_{0}>$ 0. Just as in the case of dimension one, we observe instabilities when violating the non-cavitation assumption $01+\eta_{0}>$ 0 for the sharp low-pass filter, but not the smooth low-pass filter, see Figure 6.
@@
Figure 6: Plots of the second derivative $\partial_{x}^{2}u_{N}$ of the numerical solution to (4.1) for $d=2$ at time $T=0.1$, computed with either the sharp or smooth low-pass filter. The initial data is (4.12) with $s=2,u_{{\rm l}}=0.5,v_{{\rm l}}=-0.5,u_{{\rm h}}=-v_{{\rm h}}=1$, negative minimial depth $h_{0}=-0.1$ and $N=\lfloor 2M/3\rfloor$ for $2M=2^{9}$ or $2M=2^{10}$.
@@
For the Hamiltonian Saint-Venant system (4.2), the numerical results are in line with the analysis. Whenever the hyperbolicity condition $01+\eta_{0}-(u_{0}^{2}+v_{0}^{2})>$ 0 is satisfied, the numerical scheme converges with the expected rate for both the smooth and the sharp low-pass filter. This is illustrated in Figure 5(b) for initial data (4.12) with $h_{0},u_{{\rm l}}=0.5,v_{{\rm l}}=-0.5,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$.
@@
We observe instabilities in the numerical approximation of system (4.1) when using the sharp low-pass filter for initial data violating $01+\eta_{0}-(u_{0}^{2}+v_{0}^{2})>$ 0 but not $01+\eta_{0}>$ 0. This is shown in Figure 7. There, we take as initial data (4.12) with $h_{0}=0.5,u_{{\rm l}}=2,v_{{\rm l}}=-2,s=2,u_{{\rm h}}=-v_{{\rm h}}=1$. We relate these instabilities to the lack of well-posedness of the underlying system (4.2) when the hyperbolicity condition $01+\eta_{0}-(u_{0}^{2}+v_{0}^{2})>$ 0 fails; see Lemma 4.1. The numerical scheme with the smooth low-pass filter does not exhibit instabilities for the tested values of $M$, but limited computational power prevents us from testing very large values of $M$ in dimension two.
@@
Figure 7: Plots of the second derivative $\partial_{x}^{2}u_{N}$ of the numerical solution to  (4.2) for $d=2$ at time $T=0.1$, computed with either the sharp or smooth low-pass filter. The initial data is (4.12) with $s=2,u_{{\rm l}}=2,v_{{\rm l}}=-2,u_{{\rm h}}=-v_{{\rm h}}=1$, positive minimal depth $h_{0}=0.5$ and $N=\lfloor 2M/3\rfloor$ for $2M=2^{9}$ or $2M=2^{10}$.

### v2

@@ 4.2 Numerical experiments in dimension one
where $\hat{\bm{\eta}}=(\hat{\bm{\eta}}_{-M+1},\ldots,\hat{\bm{\eta}}_{M}),\hat{\bm{u}}=(\hat{\bm{u}}_{-M+1},\ldots,\hat{\bm{u}}_{M})$ are the coefficients of the Fast Fourier transform of $\bm{\eta},\bm{u}$, and $\hat{\bm{k}}=(k/\pi\ :\ k=-M+1,\ldots,M)$ are the discrete Fourier modes. Abusing notation, we will incorrectly refer to $\hat{\bm{\eta}}_{k},\hat{\bm{u}}_{k}$ as Fourier coefficients (they are related to the coefficients $c_{j}$ of infinite Fourier series $f(x)=\sum_{j\in\mathbb{Z}}c_{j}\exp({\rm i}jx)$ through $\hat{\bm{f}}_{k}=\sum_{j\in\mathbb{Z}}c_{k+2jM}$). For functions $f\in H^{s}(2\pi\mathbb{T})$, the error due to this aliasing effect is of order ${\mathcal{O}}(M^{-s})$.
Spatial differentiation is now obtained by multiplying the Fourier coefficients with ${\rm i}k$. Nonlinear operations are computed pointwise on collocation points $x_{n}$, via inverse Fast Fourier transform. This procedure leads, in general, to aliasing errors. For polynomial nonlinearities (such as for the Saint-Venant system), one can use so-called dealiasing techniques to remove these errors. For quadratic nonlinearities, one may for example use Orszag’s $3/2$-rule [20], which consists in adding a sufficient number of Fourier modes with coefficients set to zero. For more information on spectral methods and dealiasing techniques, we refer to [5] and [25].
@@
From now on, we denote the number of collocation points by $2M$, and let $N=\lfloor 2M/3\rfloor$, that is, the greatest integer smaller than $2M/3$. Our numerical scheme maintains the highest $1/3$ of the Fourier modes to zero at each time-step, see the discussion above. Abusing notation, we will refer to the fully-discretized numerical solution as $\bm{U}_{N}=(\eta_{N},u_{N})$, since only $2N$ Fourier modes are nonzero. This convention means that $N$ plays the same role in this section as in the previous, analytical sections. We will compute the numerical solution with $2M=2^{j},j=6,\ldots,15$ collocation points and use time step $dt=10^{-5}$. The time step is an order of magnitude smaller than needed to avoid stability issues, and small enough to ensure the error due to the spatial discretization dominates. We will use the solution computed with $2M=2^{15}$ and sharp low-pass filter as a reference solution $\bm{U}_{\rm ref}=(\eta_{\rm ref},u_{\rm ref})$, and compute the relative error of the numerical solutions $\bm{U}_{N}=(\eta_{N},u_{N})$ by comparing with the reference solution:
@@
$\eta^{0}(x)=\frac{1}{2}\exp(-\left|x\right|^{\alpha})\exp(-4x^{2}),\qquad u^{0}(x)=0.$ (4.9)
Notice $\bm{U}^{0}=(\eta^{0},u^{0})$ satisfies both $1+\eta^{0}>0$ and the stricter condition $1+\eta^{0}-(u^{0})^{2}>0$. The initial surface is a heap of water situated at the origin. Both $\eta^{0},u^{0}$ decay to machine precision near $\pi,\pi$ and can therefore be seen as periodic. Moreover, $\bm{U}^{0}\in H^{\alpha+1/2}(2\pi\mathbb{T})^{2}$. We let $\alpha=1.5$ and simulate the time-evolution up to a final time $T=0.5$ with either sharp or smooth lowpass filters applied to the nonlinear terms. A plot of the initial data as well as the decay of its Fourier coefficients (through $E^{s}(\bm{U}_{N})|_{t=0}=\frac{\left|(\operatorname{Id}-{\sf P}_{N})\bm{U}_{\rm ref}|_{t=0}\right|_{H^{s}}}{\left|\bm{U}_{\rm ref}|_{t=0}\right|_{H^{s}}}$) is shown in Figure 1.
@@
Figure 2(a) show log-log plots of the error $E_{s}$ at time $T=0.5$ for the numerical solution with $2M=2^{j}$ where $j=6,\ldots,14$ computed using sharp and smooth low-pass filters. Figure 2(a) shows the relative error measured in the $L^{2}$-norm, $E_{0}(\bm{U}_{N})$ and in the $H^{1}$-norm, $E_{1}(\bm{U}_{N})$.
@@
$EOC_{s}=\frac{\log(E_{s}(\bm{U}_{N})/E_{s}(\bm{U}_{2N}))}{\log(2)},$
is given in Figure 2(b). With the initial data in $H^{2}(2\pi\mathbb{T})^{d}$, we expect from our analysis that the $L^{2}$-error should decay as ${\mathcal{O}}(N^{-2})$ and the $H^{1}$-error should decay as ${\mathcal{O}}(N^{-1})$ when using both the sharp and smooth low-pass filter. This aligns with our numerical results. The absolute error is slightly larger when using the smooth low-pass filter, which is to be expected since applying ${\sf S}_{N}$ removes more information at each time step than does ${\sf P}_{N}$. Taking as initial data (4.9) with other values of $\alpha$ (we have tested $\alpha=1,2.5,3$) also yields the expected results.
Let us now consider initial data $(\eta^{0},u^{0})\in{\mathcal{U}}\setminus{\mathcal{U}}_{\mathscr{H}}$, that is satisfying $1+\eta^{0}>0$, but not the stricter condition $1+\eta^{0}-(u^{0})^{2}>0$. In particular, we will consider the initial data from the example at the end of Section 4.1, with $q=0,p=1,s=2$:
$\bm{U}_{N}^{0}=(\eta^{0},u_{N}^{0}),\quad\eta_{N}^{0}(x)=-\frac{1}{2}\cos(x),\quad u_{N}^{0}(x)=\sin(x)+\frac{\sin(Nx)}{N^{2}}$ (4.10)
Notice $f(x)=1-\frac{1}{2}\cos(x)-\sin^{2}(x)<0$ for $x\in(-\frac{\pi}{2},-\frac{\pi}{3})\cup(\frac{\pi}{3},\frac{\pi}{2})$. On the other hand, $1+\eta^{0}(x)>0$ for all $x\in 2\pi\mathbb{T}$. We therefore expect the numerical scheme to converge when using the smooth low-pass filter, but that instabilities may emerge when discretizing with the sharp low-pass filter. Convergence plots for the numerical solutions in both cases are shown in Figure 3(b). In line with the analysis, we have convergence when using the smooth low-pass filter. However, this is also true for the sharp low-pass filter. Despite many attempts, we have not been able to observe numerical instabilities for initial data we have tested satisfying $1+\eta^{0}>0$ but violating $1+\eta^{0}-(u^{0})^{2}>0$.
@@
Interestingly, we do observe a difference between sharp and smooth low-pass filters when (barely) violating the non-cavitation assumption $1+\eta^{0}>0$. Let
@@
Figure 4 shows the second derivative of the velocity at time $T=0.1$ for smooth and sharp low-pass filters for $2M=2^{10}$ and $2M=2^{12}$. The second derivative is uniformly bounded when using the smooth low-pass filter whereas it is not around the point $x=0$ where the non-cavitation assumption is violated when using the sharp low-pass filter.
@@ 4.3 Numerical experiments in dimension two
The numerical simulations of the Saint-Venant system with $d=2$ are analogous to the case when $d=1$ and have been executed using the same Julia package WaterWaves1D [9]. They are also reproducible using the scripts available at WaterWaves1D.jl/examples/StudySaintVenant.jl. We let $2M$ denote the number of collocation points in each of the two spatial dimensions, $x$ and $y$, which form a grid with $4M^{2}$ collocation points. As in the previous section, we set $N=\lfloor 2M/3\rfloor$ and let ${\sf P}_{N},{\sf S}_{N}$ be as described in the introduction. In particular, notice ${\sf S}_{N}$ is now a composition of one-dimensional low-pass filters: ${\sf S}_{N}\coloneq\operatorname{Diag}(S_{N}(D))$ with $S_{N}((k_{1},k_{2}))\coloneq S(k_{1}/N)S(k_{2}/N)$, where we set $S(\cdot)\coloneq\max\big(0,\min\big(1,2-2|\cdot|\big)\big)^{2}$. We denote numerical approximations by $\eta_{N},u_{N},v_{N}$. The values at collocation points $\bm{\eta},\bm{u},\bm{v}$ and associated discrete Fourier modes $\bm{\hat{}}{\bm{\eta}},\hat{\bm{u}},\hat{\bm{v}}$ are $2M$-by-$2M$ matrices.
Our main interest in studying numerically the systems in two dimensions is to examine whether we observe any difference between the Hamiltonian and non-Hamiltonian version of the Saint-Venant system, respectively (4.2) and (4.1). While there is indeed a difference with respect to stability of the numerical schemes —we observe instabilities with the sharp low-pass filter when violating $1+\eta^{0}-(u^{0})^{2}-(v^{0})^{2}>0$ for (4.2), but not for (4.1)— this can be explained by the difference between the hyperbolicity domains $\mathcal{U}$$_{\mathscr{H}}$ and $\mathcal{U}$} rather than by the presence or absence of a Hamiltonian structure. Because the setting of dimension $d=2$ is computationally costlier, we use timestep $dt=5\times 10^{-4}$ in our numerical experiments and, when calculating relative errors, we take as a reference solution the numerical solution computed with $2M=2^{10}$.
@@
$\begin{split}\eta^{0}(x,y)&=(h_{0}-1)\cos(x)\cos(y),\\ u^{0}(x,y)&=u_{{\rm l}}\sin(x)\cos(y)+u_{{\rm h}}\frac{\sin(Nx)\cos(Ny)}{N^{s}},\\ v^{0}(x,y)&=v_{{\rm l}}\cos(x)\sin(y)+v_{{\rm h}}\frac{\cos(Nx)\sin(Ny)}{N^{s}},\end{split}$ (4.12)
@@
For the standard, non-Hamiltonian Saint-Venant system (4.1) our numerical results when $d=2$ align with the numerical results when $d=1$. That is, the numerical approximation converges with order $s$ for tested initial data in $H^{s}(2\pi\mathbb{T})$ as long as $1+\eta^{0}>0$ when using both the smooth and sharp low-pass filters. Figure 5(a) shows a log-log plot of the relative error measured in the $L^{2}$-norm, $E_{0}(\bm{U}_{N})$ and in the $H^{1}$-norm, $E_{1}(\bm{U}_{N})$, at time $T=0.1$ for $2M=2^{j},j=5,\ldots,9$ and $N=\lfloor 2M/3\rfloor$ with initial data (4.12) with $h_{0}=0.5,u_{{\rm l}}=-v_{{\rm l}}=0.5,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$. Analogous results also hold for other values of $s$ (we have tested $s=2.5,3$).
@@
Figure 5: Plot illustrating the convergence of the numerical schemes (4.8) and (4.7) for the systems (4.1) (in the left) and (4.2) (in the right) in two spatial dimensions as the number of collocation points $2M$ increases. The plot shows the relative error of the numerical solution for initial data (4.12) with $h_{0}=0.5,u_{{\rm l}}=-v_{{\rm l}}=0.5,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$ at time $T=0.1$. The initial data is in in $H^{2}((2\pi\mathbb{T})^{2})^{3}$ and the relative error is measured in the $L^{2}$-norm, $E_{0}$ and in the $H^{1}$-norm, $E_{1}$ for $2M=2^{j},j=6,\ldots,9$ when using either sharp or smooth low-pass filters. To illustrate, the blue and orange lines have slopes 2-2 and 1-1 respectively. The numerical scheme exhibits spectral convergence with both sharp and smooth low-pass filters, for both systems.
@@
We have not observed any instabilities in the numerical approximation of system (4.1) due to the use of sharp low-pass filter for any of the initial data we tested satisfying $1+\eta^{0}>0$. Just as in the case of dimension one, we observe instabilities when violating the non-cavitation assumption $1+\eta^{0}>0$ for the sharp low-pass filter, but not the smooth low-pass filter, see Figure 6.
@@
Figure 6: Plots of the second derivative $\partial_{x}^{2}u_{N}$ of the numerical solution to (4.1) for $d=2$ at time $T=0.1$, computed with either the sharp or smooth low-pass filter. The initial data is (4.12) with $s=2,u_{{\rm l}}=-v_{{\rm l}}=0.5,u_{{\rm h}}=-v_{{\rm h}}=1$, negative minimial depth $h_{0}=-0.1$ and $N=\lfloor 2M/3\rfloor$ for $2M=2^{9}$ or $2M=2^{10}$.
@@
For the Hamiltonian Saint-Venant system (4.2), the numerical results are in line with the analysis. Whenever the hyperbolicity condition $1+\eta^{0}-(u^{0})^{2}-(v^{0})^{2}>0$ is satisfied, the numerical scheme converges with the expected rate for both the smooth and the sharp low-pass filter. This is illustrated in Figure 5(b) for initial data (4.12) with $h_{0},u_{{\rm l}}=-v_{{\rm l}}=0.5,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$.
@@
We observe instabilities in the numerical approximation of system (4.2) when using the sharp low-pass filter for initial data violating $1+\eta^{0}-(u^{0})^{2}-(v^{0})^{2}>0$ but not $1+\eta^{0}>0$. This is shown in Figure 7. There, we take as initial data (4.12) with $h_{0}=0.5,u_{{\rm l}}=-v_{{\rm l}}=2,u_{{\rm h}}=-v_{{\rm h}}=1$ and $s=2$. We relate these instabilities to the lack of well-posedness of the underlying system (4.2) when the hyperbolicity condition $1+\eta^{0}-(u^{0})^{2}-(v^{0})^{2}>0$ fails; see Lemma 4.1. The numerical scheme with the smooth low-pass filter does not exhibit instabilities for the tested values of $M$, but limited computational power prevents us from testing very large values of $M$ in dimension two.
@@
Figure 7: Plots of the second derivative $\partial_{x}^{2}u_{N}$ of the numerical solution to (4.2) for $d=2$ at time $T=0.1$, computed with either the sharp or smooth low-pass filter. The initial data is (4.12) with $s=2,u_{{\rm l}}=-v_{{\rm l}}=2,u_{{\rm h}}=-v_{{\rm h}}=1$, positive minimal depth $h_{0}=0.5$ and $N=\lfloor 2M/3\rfloor$ for $2M=2^{9}$ or $2M=2^{10}$.